Homework: Homework 2.1 Score: 0 of 1 p HW Score: 64.76%, 9 07 Question He Students in an experimental psychology class did research on depression as a sign of stress. A test was administered to a sample of 30 students. The scores 6 of 14 (12 complete 2.365-T given. Complete parts (a) through (c) below 29 ST 47 47 56 24 (a) Find the mean and median of the data The mean is (Round to three decimal places as needed) Enter your answer in the answer box and then cick Check Answer 3 Cear A here to search 5
Answer #1

Solution:

Given that,

a ) The sample mean is \bar x

Mean \bar x  = \sum/ n

=  (29+53+46+39+47+52+59+38+46+47+57+18+34+56+24+42+26+32+44+39+54+45+42+59+38+45+24+58+38+13)/30

= 124430

= 41.4667

Mean \bar x = 41.467

b ) Median :
Observations in the ascending order are :
13,18,24,24,26,29,32,34,38,38,38,39,39,42,42,44,45,45,46,46,47,47,52,53,54,56,57,58,59,59

n = 30

M = ( Value of ( n / 2 )th  Observations +  Value of (( n / 2 ) + 1 )th  Observations ) / 2

= ( Value of ( 30 / 2 )th  Observations +  Value of ((30 / 2 ) + 1 )th  Observations ) / 2

= ( Value of 15th  Observations +  Value of 16th  Observations ) / 2

= ( 42 + 44 ) / 2

= 86 / 2

= 43

Median = 43